OpenStudy (danmac0710):

f(x) is a relation on a real variable x such that: e^f(x) - f(x) = e^5 - 5..... find value of y coordinate and the gradient when x=1. Use Implicit differentiation. Thanks for any light you can shed on this one!

OpenStudy (anonymous):

\[e^y-y=e^5-5\] like that?

OpenStudy (anonymous):

because the right hand side is a constant, so if that is the question, then \[e^yy'-y'=0\]

OpenStudy (anonymous):

well that didn't help did it? i guess i am stumped

OpenStudy (danmac0710):

Thanks for trying though Satellite73!